OpenStudy (anonymous):

The number of days needed to finish a quilt varies inversely as the number of people working on it. If it takes 15 days for three people to finish a quilt, how many people would be needed to finish it in 2.5 days? 0.5 18 12 20

OpenStudy (anonymous):

y - days x - people y=k/x 15=k/3 => k=45 so equation is y =45/x 2,5=45/x => x=45/2,5 =18